OpenStudy (anonymous):

How do you determine the range of a radical function such as f(x) = 3- sqrt (2x+5)

OpenStudy (across):

Well, if f is a function mapped to the real numbers, x cannot be less than 5/2, which would leave us with a range as follows: \[3 \le f(x) \le \infty\]

OpenStudy (across):

Pardon, it should be strictly less than infinity. ;p

OpenStudy (anonymous):

Got it. Thanks very much.